OpenStudy (anonymous):

if we are integrating trig functions in cubic or squared form , do we have to convert them into linear form first? And can anyone tell me the steps of differential equations? Thanks

OpenStudy (anonymous):

Like using identities to convert them into linear form

OpenStudy (callisto):

Usually, yes, but it depends For example, for \[\int cos^2x dx\]\[=\int \frac{cos2x+1}{2} dx\]\[=\frac{1}{2} \int cos2x+1 dx = \frac{1}{4}sin2x + \frac{x}{2}+C\]

OpenStudy (callisto):

But when you're integrating \[\int sin^3xcos^3x dx\] That's another story \[\int sin^3xcos^3x dx\]\[=\int sin^3xcos^2xcosx dx\]\[=\int sin^3x(1-sin^2x) d(sinx)\]\[=\int sin^3x-sin^5x d(sinx)\]\[=\frac{sin^4x}{4}-\frac{sin^6x}{6} +C\]